OpenStudy (anonymous):

How many multiples of 3 are there between 100 and 1000?

OpenStudy (anonymous):

@ranga @SolomonZelman

OpenStudy (anonymous):

300

OpenStudy (anonymous):

how did you get that answer though?

OpenStudy (anonymous):

1000/3 = 333.33 Therefor there are 333 numbers between 0 and 1000 100/3 = 33.333 Therefor there are 33 numbers between 0 and 100 Betwen 100 and 1000 333 - 33 = 300 There are 300 numbers divisible by 3 between 100 and 1000

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks for explaining @Mason_Jones573